OpenStudy (anonymous):

A line passes through the origin and has a slope of 4. What is the equation of the line that is parallel to the first line and passes through the point (1, –2)?

OpenStudy (anonymous):

y=mx+b , m=4 (slope) Lets find the B of the line that passes through the point (1, -2) -2=4*(1)+b b=-6 Therefore, with a line that is parallel, the slope will stay the same, only the B value will change since that is your vertical intercept value. y=4x-6 is the equation of the line that is parallel to a line with a slope of 4 that passes through 0 , 0

OpenStudy (anonymous):

another way to do this is using point slope form y-y1=m(x-x1) m=4 and (x1,y1)=(1,-2) y+2=4(x-1) y+2=4x-4 y=4x-6